OpenStudy (anonymous):

If t1 = 5 and tn = tn-1 + 2, find the first 4 terms of the sequence.

OpenStudy (anonymous):

\[t_2=t_1+2\] so \[t_3=t_2+2\] and \[t_4=t_3+2\] just calculate t\2 cuz you know t_1.....

OpenStudy (anonymous):

IAM CONFUSED

OpenStudy (anonymous):

I'll do the first one kk ?!

OpenStudy (anonymous):

to compute t_2 I'll take n=2 an see what I'll get \[t_2=t_{2-1}+2=t_1+2\] as I know t_1=1 so t_2=1+2=3 did you understand .?

OpenStudy (anonymous):

sorry t_1=5 ! so t_2=5+2=7

OpenStudy (anonymous):

can you do the same for the rest ?!

OpenStudy (anonymous):

I THINK SO

OpenStudy (anonymous):

so do it ! I can check ! kk ?

OpenStudy (anonymous):

IM SORRY I STILL DONT UNDERSTAND I THINK I REALLY DONT UNDERSTAND WHAT THE QUESTION IS ASKING ME WHAT AM I SUPPOSED TO DO EXACTLY LIKE WHAT IS THE QUESTION ASKING ME

OpenStudy (anonymous):

THis is something called a recursive function. Plug in values for n. \[t_4 = t_3 + 2\]\[t_3 = t_2 + 2\]\[t_2 = t_1 + 2\]\[t_1 = 5\]You have \(t_1\) = 5, so just plug that into the above function. \[t_2 = t_1 + 2 \implies t_2 = 5 + 2 \implies t_2 = 7\]Can you do the same for the rest that's above?

OpenStudy (anonymous):

SO IM FINDING WHAT THE OTHERS MEAN

OpenStudy (anonymous):

Yes. If you realize, each of the functions include the value from before. So if \(t_2\) = 7, then plug it into \(t_3 = t_2 + 2\)and do the same again using \(t_3\) and you'll have you're answer :)

OpenStudy (anonymous):

SO THAT WILL GIVE ME 4 TERMS RIGHT

OpenStudy (anonymous):

Yes. 5 is one and 7 is another one like I said. Can you figure out the other 2?

OpenStudy (anonymous):

YES THANK YOU VERY MUCH YU EXPLAINED IT SO I WOULD UNDERSTAND IT

OpenStudy (anonymous):

np :)